OpenStudy (nightmarenight):

Help??!!! Will fan and give medal for correct answer! Evaluate the function T(x) = 19 – x, when x = –3

OpenStudy (anonymous):

Replace the x in the equation with the value of x given, which is -3: T(-3)=19-(-3) T(-3)=19+3 T(-3)=22